OpenStudy (anonymous):

The length of a rectangle is 12 m and its width is 3 m. Find the area of ¼ of the rectangle.

OpenStudy (mertsj):

What is the area of the whole rectangle?

OpenStudy (anonymous):

wouldn't it be (1/4)12 * 3

OpenStudy (anonymous):

A = B * H

OpenStudy (anonymous):

12*3=36/4=9